OpenStudy (anonymous):

Will give medal! Explain what is accomplished by factoring. As part of your explanation, write a sample problem in which you factor a monomial from a polynomial. Then write and factor another problem that is a quadratic expression. Check your solutions by multiplying after you factor. Thanks.

OpenStudy (***[isuru]***):

factoring is basically done to ease the process of simplification ......... if u given \[\frac{ ( x^{2} -4x + 4) }{ ( x - 2)^{2} }\] and ask u to simplify it .... u will not able to think of anything... but if u factorize it u will get \[\frac{ (x -2)^{2} }{ (x -2)^{2} } = 1\] See.... u was able to do that because u were able to factorize the denominator.. i f u couldnt do that.. u would never simplify that... and there r many more uses... this one is pretty basic hope this will help ya :)

OpenStudy (anonymous):

Thanks (:

OpenStudy (***[isuru]***):

u r welcome !